Suppose now the government decides to increase the number of quotas available to 72 units, but it keeps the price support at the current level of $72,

calculate

the consumer surplus

the producer surplus

deadweight loss


1
Expert's answer
2021-10-15T11:23:24-0400

Solution:

Consumer surplus = "\frac{1}{2} \times(72-50) (72-45) = \frac{1}{2} \times(22 \times27) =\frac{1}{2} \times594 = 297"


Producer surplus = "\frac{1}{2} \times(72-68) (65-30) = \frac{1}{2} \times(4 \times35) =\frac{1}{2} \times140 = 70"


Deadweight loss = "\frac{1}{2} \times(45-38) (45-30) + \frac{1}{2} \times(72-68) (45-30)= \frac{1}{2} \times(7 \times15) + \frac{1}{2} \times(4 \times15) =\frac{1}{2} \times105 + \frac{1}{2} \times60 = 53 + 30 = 83"